# Stability Analysis of Thermohaline Convection With a Time-Varying Shear Flow Using the Lyapunov Method

### v1

@@ Abstract
This work identifies instabilities and computes the growth rate of a linear time-varying system using the Lyapunov method. The linear system describes cold fresh water on top of hot salty water with a periodically time-varying background shear flow. We employ a time-dependent weighting matrix to construct a Lyapunov function candidate, and the resulting linear matrix inequalities formulation is discretized in time using the forward Euler method. As the number of temporal discretization points increases, the growth rate predicted from the Lyapunov method or the Floquet theory will converge to the same value as that obtained from numerical simulations. We also use the Lyapunov method to analyze the instantaneous principal direction of instabilities and compare the computational resources required by the Lyapunov method, numerical simulations, and the Floquet theory.
@@ I Introduction
Stability analysis can provide key insights into pattern formation and the transition to chaotic and turbulent behaviors [6], but the time dependence of background flow introduces additional challenges in stability analysis [7]. Quasi-steady assumption and eigenvalue analysis of the time-frozen base state are known to oversimplify the dynamics of a time-varying system, and thereby may miss instabilities induced by the time-dependence [8, 9]. Floquet theory [10] is a powerful tool to analyze the stability of a periodic orbit, but it is limited to the linear time-periodic system [7]. Stability of time-varying systems can also be analyzed by numerical simulations [4, 5]. However, simulations only provide realizations of certain initial conditions to support stability analysis, but cannot formally prove the stability. For nonlinear systems, extensive simulations with random initial conditions (e.g., 100,000 simulations [11]) might be needed to fully characterize the region of attraction.
@@
The Lyapunov method has been widely used to analyze the stability of nonlinear systems, and can be generalized to investigate the stability of time-varying (non-autonomous) systems [8, 12]. A periodic Lyapunov function is introduced to analyze a discrete-time nonlinear periodically time-varying system, which was shown to provide a larger region of attraction estimation compared with standard Lyapunov theory [13]. Moreover, stabilization of continuous-time periodic linear systems can be achieved by solving a periodic Lyapunov differential equation [14]. The Lyapunov method can also be applied to analyze the nonlinear stability (e.g., characterization of region of attraction [15]) or analyze non-periodically time-varying systems [16], which can overcome limitations of numerical simulations and Floquet theory.
@@
This work will employ the Lyapunov method to analyze instabilities of a linear time-periodic system that describes thermohaline-shear instability [4], where cold fresh water on top of hot salty water interacts with a periodically time-varying shear flow. We will use a periodic weighting matrix $\boldsymbol{P}(t)$ to construct a Lyapunov function candidate $V(\boldsymbol{\psi},t)=\boldsymbol{\psi}^{*}\boldsymbol{P}(t)\boldsymbol{\psi}$, where $\boldsymbol{\psi}$ is the state variable, and the superscript $(\cdot)^{*}$ means the Hermitian of the argument. We formulate linear matrix inequalities (LMI) to obtain an upper bound of the growth rate, and the LMI is discretized in time by a forward Euler method. The resulting LMI will be coupled over time and solve the time-varying weighting matrix $\boldsymbol{P}(t)$ over one period. As the number of temporal discretization points increases, the growth rate predicted from the Lyapunov method or the Floquet theory converges to that obtained from numerical simulations. We also perform an eigendecomposition of the instantaneous weighting matrix $\boldsymbol{P}(t)$ to uncover the principal direction of thermohaline-shear instability. The computational resources requirement is also compared among the Lyapunov method, numerical simulations, and the Floquet theory.
@@
The paper is organized as follows. Section II introduces the linear time-varying system that describes thermohaline-shear instabilities [4]. Section III presents three methods that will be used to compute growth rate, including the Lyapunov method, numerical simulations, and the Floquet theory. Section IV presents the results, and we conclude this paper with a discussion of future directions in Section V.
@@ II Problem Formulation
We consider the flow configuration in Fig. 1, where we have a time-varying background shear flow $\overline{\boldsymbol{U}}_{*}(z_{*},t_{*})=[A_{U*}(t_{*})z_{*},A_{V*}(t_{*})z_{*},0]^{\text{T}}$. We consider cold fresh water on top of hot salty water with a large-scale background temperature gradient $\partial_{z_{*}}\overline{T}_{*}<0$ and salinity gradient $\partial_{z_{*}}\overline{S}_{*}<0$. We assume that the density is influenced by temperature and salinity linearly $(\rho_{*}-\rho_{0*})/\rho_{0*}=\beta(S_{*}-S_{0*})-\alpha(T_{*}-T_{0*})$, where $\alpha$ and $\beta$ are expansion and contraction coefficients. Here, $\rho_{0*}$, $T_{0*}$, and $S_{0*}$ are reference density, temperature, and salinity, respectively. We define dimensionless length, time, velocity, and pressure as $z=z_{*}/d$, $t=t_{*}/(d^{2}/\kappa_{T})$, $u=u_{*}/(\kappa_{T}/d)$, and $p=p_{*}/(\rho_{0*}\nu\kappa_{T}/d^{2})$ [4], where $d:=\left[\kappa_{T}\nu/(g\alpha|\partial_{z_{*}}\overline{T}_{*}|)\right]^{1/4}$, $\kappa_{T}$ is the thermal diffusivity, and $\nu$ is the viscosity. We normalize temperature and salinity by $T=T_{*}/(d\,|\partial_{z_{*}}\overline{T}_{*}|)$ and $S=\beta S_{*}/(\alpha\,d\,|\partial_{z_{*}}\overline{T}_{*}|)$, and then we have dimensionless background temperature as $\overline{T}=-z$ and background salinity as $\overline{S}=-R_{\rho}z$. After linearization around the base state $(\overline{T},\overline{S},\overline{\boldsymbol{U}})$, we have the non-dimensional governing equation of fluctuations $(T,S,\boldsymbol{u})$ [4] as:
@@
Here, we have the density ratio $R_{\rho}:=\beta\partial_{z_{*}}\overline{S}_{*}/(\alpha\partial_{z_{*}}\overline{T}_{*})$, the Prandtl number $Pr:=\nu/\kappa_{T}$, and the diffusivity ratio $\tau:=\kappa_{S}/\kappa_{T}$, where $\nu$ is the viscosity and $\kappa_{S}$ is the salinity diffusivity. The unit vector in the vertical $z$ direction is denoted as $\boldsymbol{e}_{z}$. Here, we consider the horizontal direction $(x,y)$ as periodic. The dimensionless background shear flow is $\overline{\boldsymbol{U}}(z,t)=[\overline{u}(z,t),\overline{v}(z,t),0]^{\text{T}}=[A_{U}(t)z,A_{V}(t)z,0]^{\text{T}}$ with an unbounded vertical direction $z$, i.e., unbounded Couette flow model. Thus, we can perform the following solution ansatz [4] to eliminate the dependence on $z$:
@@
where the state variable $\boldsymbol{\psi}=\begin{bmatrix}\hat{T}&\hat{S}&\hat{u}&\hat{v}&\hat{w}\end{bmatrix}^{\text{T}}$. The matrix $\boldsymbol{A}(t)$ is defined in (LABEL:eq:a_matrix) with $c(t)=k^{2}+l^{2}+m(t)^{2}$:
@@ III Methods
In this section, we present our formulation to compute the growth rate of the linear time-varying system in (5) using the Lyapunov method. We then show how to obtain the growth rate using numerical simulations and the Floquet theory, which will be used as a basis for the comparison with the Lyapunov method.
@@ III-A Lyapunov stability analysis
Given a linear time-varying system $\dot{\boldsymbol{\psi}}=\boldsymbol{A}(t)\boldsymbol{\psi}$ in (5). If we can find a continuously differentiable Hermitian matrix $\boldsymbol{P}(t)\in\mathbb{C}^{5\times 5}$ and $\overline{\lambda}$ by solving
@@
$\;\;\;\;\dot{\boldsymbol{P}}(t)+\boldsymbol{A}^{*}(t)\boldsymbol{P}(t)+\boldsymbol{P}(t)\boldsymbol{A}(t)-2\overline{\lambda}\boldsymbol{P}(t)\preceq 0,$ (7c)
then $\|\boldsymbol{\psi}(t)\|_{2}\leq Ce^{\overline{\lambda}t}\|\boldsymbol{\psi}(0)\|_{2}$, where $C$ is a constant.
@@
We consider a function $V(\boldsymbol{\psi},t)=\boldsymbol{\psi}^{*}\boldsymbol{P}(t)\boldsymbol{\psi}$ and compute its time derivative $\dot{V}$ as:
@@
$\leq 2\overline{\lambda}\boldsymbol{\psi}^{*}\boldsymbol{P}(t)\boldsymbol{\psi}=2\overline{\lambda}\,V,$ (8c)
where (8c) is obtained by the inequality in (7c). Integrating the inequality (8) over time $t$, we have
$V(\boldsymbol{\psi},t)\leq e^{2\overline{\lambda}t}V(\boldsymbol{\psi},0).$ (9)
We then use the Rayleigh-Ritz theorem to obtain
@@
where $\mu_{\text{min}}[\cdot]$ and $\mu_{\text{max}}[\cdot]$ represent the minimum and maximum eigenvalues of the argument. Combining the bound on $V(\boldsymbol{\psi},t)$ in (9) with (10) evaluated at $t=0$, we obtain
$\mu_{\text{min}}[\boldsymbol{P}(t)]$ $\|\boldsymbol{\psi}(t)\|_{2}^{2}\leq V(\boldsymbol{\psi},t)$ (11a)
$\leq$ $e^{2\overline{\lambda}t}V(\boldsymbol{\psi},0)\leq e^{2\overline{\lambda}t}\mu_{\text{max}}[\boldsymbol{P}(0)]\|\boldsymbol{\psi}(0)\|_{2}^{2}.$ (11b)
Thus, we have
$\|\boldsymbol{\psi}(t)\|_{2}$ $\leq\sqrt{\frac{\mu_{\text{max}}[\boldsymbol{P}(0)]}{\mu_{\text{min}}[\boldsymbol{P}(t)]}}e^{\overline{\lambda}t}\|\boldsymbol{\psi}(0)\|_{2}$ (12a)
$\leq\sqrt{\frac{\mu_{\text{max}}[\boldsymbol{P}(0)]}{\epsilon}}e^{\overline{\lambda}t}\|\boldsymbol{\psi}(0)\|_{2}.$ (12b)
Here, the inequality in (12b) is obtained by (7b). ∎
We then discretize $\boldsymbol{P}(t)$ as $\boldsymbol{P}(t_{i})$ and discretize $\boldsymbol{A}(t)$ as $\boldsymbol{A}(t_{i})$, i=0,1,,ni=0,1,...,n with a fixed time step $\Delta t=t_{i+1}-t_{i}$. The forward Euler method is used to approximate the time derivative as $\dot{\boldsymbol{P}}(t_{i})=\frac{\boldsymbol{P}(t_{i+1})-\boldsymbol{P}(t_{i})}{\Delta t}$, which converts the LMI in (7) as:
@@
$\dot{\boldsymbol{P}}(t_{i})+\boldsymbol{A}^{*}(t_{i})\boldsymbol{P}(t_{i})+\boldsymbol{P}(t_{i})\boldsymbol{A}(t_{i})-2\overline{\lambda}\boldsymbol{P}(t_{i})\preceq 0.$ (13c)
For a time-periodic system, we can enforce periodicity on $\boldsymbol{P}(t)$ such that $\boldsymbol{P}(t_{n})=\boldsymbol{P}(t_{0})$ [13, 14], and we only need to consider one period for the discretized LMI. Thus, we select $t_{0}=0$ and $t_{n}=T$, where $T=\frac{2\pi}{\omega}$ is the oscillation period. We also found that using a time-invariant $\boldsymbol{P}$ to construct a Lyapunov function candidate will result in the growth rate being overly conservative.
Because $\boldsymbol{A}(t)$ in (LABEL:eq:a_matrix) is a real matrix, we can assume the weighting matrix $\boldsymbol{P}(t)\in\mathbb{R}^{5\times 5}$ and $\boldsymbol{P}(t_{i})\in\mathbb{R}^{5\times 5}$ as real symmetric matrices without loss of generality. The value of $\epsilon$ does not influence the upper bound of the growth rate $\overline{\lambda}$ due to the homogeneity of the constraints in (7c) and (13c), and thus, we set $\epsilon=1$ without loss of generality. We use YALMIP [17] to formulate our LMI in (13) and use MOSEK [18] as the semi-definite programming (SDP) solver. We first check the stability of the system by checking the feasibility of constraints in (13) with $\overline{\lambda}=0$, which is much quicker than calculating the value of $\overline{\lambda}$. If infeasible, we then solve (13) to minimize the upper bound of the growth rate $\overline{\lambda}>0$ through bisection search over $\overline{\lambda}$.

### v2

@@ Abstract
This work demonstrates that the Lyapunov method can effectively identify the growth rate of a linear time-periodic system describing cold fresh water on top of hot salty water with a periodically time-varying background shear flow. We employ a time-dependent weighting matrix to construct a Lyapunov function candidate, and the resulting linear matrix inequalities are discretized in time using the forward Euler method. As the number of temporal discretization points increases, the growth rate predicted from the Lyapunov method or the Floquet theory will converge to the same value as that obtained from numerical simulations. Additionally, the Lyapunov method is used to analyze the most dangerous disturbance, and we also compare computational resource usage for the Lyapunov method, numerical simulations, and the Floquet theory.
@@ I Introduction
Stability analysis can provide key insights into pattern formation and the transition to chaotic and turbulent behaviors [7], but the time dependence of background flow introduces additional challenges in stability analysis [5]. Quasi-steady assumption and eigenvalue analysis of the time-frozen base state are known to oversimplify the dynamics of a time-varying system, and thereby may miss instabilities induced by the time-dependence [10, 11]. Floquet theory [6] is a powerful tool to analyze the stability of a periodic orbit, but Floquet methods are not applicable for nonlinear, non-periodic systems [5]. Stability of time-varying systems can also be analyzed by numerical simulations [17, 16]. However, simulations only support stability analysis through realizations of certain initial conditions, requiring exhaustive random search for nonlinear systems. For example, extensive simulations with more than 100,000 random initial conditions were used to characterize the region of attraction of a nine-mode shear flow model [9].
@@
Unlike numerical simulation, Lyapunov-based stability analysis does not require extensive simulations, but directly characterizes the solution trajectory for a set of initial conditions. The Lyapunov method has also been generalized to investigate the stability of time-varying (non-autonomous) systems [10, 3]. For example, a periodic Lyapunov function is introduced to analyze a discrete-time nonlinear periodically time-varying system [2], and stabilization of continuous-time periodic linear systems can be achieved by solving a periodic Lyapunov differential equation [22]. The Lyapunov method can also be applied to analyze nonlinear stability [13], nonlinear input-output properties [20], and non-periodically time-varying systems [21], which can overcome limitations of Floquet theory. Here, we will analyze the special case of a linear periodically time-varying system describing oceanic flows using the Lyapunov method, with the potential to extend to more general systems in mind. We demonstrate that the Lyapunov method can characterize the growth rate consistent with Floquet theory and numerical simulations for the defined system.
@@
Our model describes thermohaline-shear instability [17] where cold fresh water rests on top of hot salty water that interacts with a periodically time-varying shear flow. We will use a periodic weighting matrix $\boldsymbol{P}(t)$ to construct a Lyapunov function candidate $V(\boldsymbol{\psi},t)=\boldsymbol{\psi}^{*}\boldsymbol{P}(t)\boldsymbol{\psi}$, where $\boldsymbol{\psi}$ is the state variable, and the superscript $(\cdot)^{*}$ means the Hermitian of the argument. We formulate linear matrix inequalities (LMI) to obtain an upper bound of the growth rate, and the LMI is discretized in time by a forward Euler method. The resulting LMI will be coupled over time to solve the time-varying weighting matrix $\boldsymbol{P}(t)$ over one period. As the number of temporal discretization points increases, the growth rate predicted from the Lyapunov method or the Floquet theory converges to that obtained from numerical simulations. We also perform an eigendecomposition of the instantaneous weighting matrix $\boldsymbol{P}(t)$ to uncover the most dangerous disturbance of thermohaline-shear instability. Computational resource usage is compared among the Lyapunov method, numerical simulations, and Floquet theory.
@@
The paper is organized as follows. Sec. II introduces the linear time-varying system that describes thermohaline-shear instabilities [17]. Sec. III presents how to compute growth rate using the Lyapunov method, numerical simulations, and the Floquet theory. Sec. IV presents the results, and Sec. V concludes this paper with a discussion of future directions.
@@ II Problem Formulation
We consider the flow configuration in Fig. 1, where we have a time-varying background shear flow $\overline{\boldsymbol{U}}_{*}(z_{*},t_{*})=[A_{U*}(t_{*})z_{*},A_{V*}(t_{*})z_{*},0]^{\text{T}}$. We consider cold fresh water on top of hot salty water with a large-scale background temperature gradient $\partial_{z_{*}}\overline{T}_{*}<0$ and salinity gradient $\partial_{z_{*}}\overline{S}_{*}<0$. We assume that the density is influenced by temperature and salinity linearly $(\rho_{*}-\rho_{0*})/\rho_{0*}=\beta(S_{*}-S_{0*})-\alpha(T_{*}-T_{0*})$, where $\alpha$ and $\beta$ are expansion and contraction coefficients. Here, $\rho_{0*}$, $T_{0*}$, and $S_{0*}$ are reference density, temperature, and salinity, respectively. We define dimensionless length, time, velocity, and pressure as $z=z_{*}/d$, $t=t_{*}/(d^{2}/\kappa_{T})$, $u=u_{*}/(\kappa_{T}/d)$, and $p=p_{*}/(\rho_{0*}\nu\kappa_{T}/d^{2})$ [17], where $d:=\left[\kappa_{T}\nu/(g\alpha|\partial_{z_{*}}\overline{T}_{*}|)\right]^{1/4}$, $\kappa_{T}$ is the thermal diffusivity, and $\nu$ is the viscosity. We normalize temperature and salinity by $T=T_{*}/(d\,|\partial_{z_{*}}\overline{T}_{*}|)$ and $S=\beta S_{*}/(\alpha\,d\,|\partial_{z_{*}}\overline{T}_{*}|)$. Then, we have dimensionless background temperature as $\overline{T}=-z$ and background salinity as $\overline{S}=-R_{\rho}z$ to model the transition from hot salty water on the bottom to cold fresh water on the top (Fig. 1). After linearization around the base state $(\overline{T},\overline{S},\overline{\boldsymbol{U}})$, we have the non-dimensional governing equation of fluctuations $(T,S,\boldsymbol{u})$ [17] as:
@@
Here, we have the density ratio $R_{\rho}:=\beta\partial_{z_{*}}\overline{S}_{*}/(\alpha\partial_{z_{*}}\overline{T}_{*})$, the Prandtl number $Pr:=\nu/\kappa_{T}$, and the diffusivity ratio $\tau:=\kappa_{S}/\kappa_{T}$, where $\nu$ is the viscosity and $\kappa_{S}$ is the salinity diffusivity. The unit vector in the vertical $z$ direction is denoted as $\boldsymbol{e}_{z}$. The dimensionless background shear flow is $\overline{\boldsymbol{U}}(z,t)=[\overline{u}(z,t),\overline{v}(z,t),0]^{\text{T}}=[A_{U}(t)z,A_{V}(t)z,0]^{\text{T}}$ with an unbounded vertical direction $z$, allowing us to consider the spatial domain $x,y,z\in(-\infty,\infty)$ with periodic boundary conditions for fluctuations $(T,S,\boldsymbol{u})$ in $x$, $y$, and $z$ directions. Thus, we can perform the following solution ansatz [17]:
@@
where the state variable $\boldsymbol{\psi}=\begin{bmatrix}\hat{T}&\hat{S}&\hat{u}&\hat{v}&\hat{w}\end{bmatrix}^{\text{T}}\in\mathbb{C}^{5\times 1}$ are the Fourier coefficients. The matrix $\boldsymbol{A}(t)\in\mathbb{R}^{5\times 5}$ is defined in (LABEL:eq:a_matrix) with $c(t)=k^{2}+l^{2}+m(t)^{2}$:
The detailed derivation of Eq. (LABEL:eq:a_matrix) is available at Ref. [12].
@@ III Methods
In this section, we present our formulation to compute the growth rate of the linear time-varying system in (5) using the Lyapunov method. We then show how to obtain the growth rate using numerical simulations and the Floquet theory, which will be used as a basis for comparison with the Lyapunov method.
@@ III-A Lyapunov stability analysis
Given a linear time-varying system $\dot{\boldsymbol{\psi}}=\boldsymbol{A}(t)\boldsymbol{\psi}$ with $\boldsymbol{\psi}\in\mathbb{C}^{5\times 1}$. If we can find a continuously differentiable Hermitian matrix $\boldsymbol{P}(t)\in\mathbb{C}^{5\times 5}$ and $\overline{\lambda}$ by
@@
$\;\;\;\;\dot{\boldsymbol{P}}(t)+\boldsymbol{A}(t){\color[rgb]{0,0,0}{}^{*}}\boldsymbol{P}(t)+\boldsymbol{P}(t)\boldsymbol{A}(t)-2\overline{\lambda}\boldsymbol{P}(t)\preceq 0,$ (7c)
then $\|\boldsymbol{\psi}(t)\|_{2}\leq Ce^{\overline{\lambda}(t-t_{0})}\|\boldsymbol{\psi}(t_{0})\|_{2}$, where $C$ is a constant.
@@
Consider a function $V(\boldsymbol{\psi},t)=\boldsymbol{\psi}^{*}\boldsymbol{P}(t)\boldsymbol{\psi}$ and compute its time derivative $\dot{V}$ as:
@@
$\leq 2\overline{\lambda}\,\boldsymbol{\psi}^{*}\boldsymbol{P}(t)\boldsymbol{\psi}=2\overline{\lambda}\,V.$ (8c)
Thus, we have $V(\boldsymbol{\psi}(t),t)\leq e^{2\overline{\lambda}(t-t_{0})}V(\boldsymbol{\psi}(t_{0}),t_{0})$. We then use the Rayleigh-Ritz theorem [8, Theorem 4.2.2] to obtain
@@
where $\mu_{\text{min}}[\cdot]$ and $\mu_{\text{max}}[\cdot]$ represent the minimum and maximum eigenvalues of the argument.
Combining the bound on $V(\boldsymbol{\psi},t)$ and Eq. (9), we obtain
$\mu_{\text{min}}[\boldsymbol{P}(t)]\|\boldsymbol{\psi}(t)\|_{2}^{2}\leq V(\boldsymbol{\psi}(t),t)\leq e^{2\overline{\lambda}(t-t_{0})}V(\boldsymbol{\psi}(t_{0}),t_{0})$
$\leq e^{2\overline{\lambda}(t-t_{0})}\mu_{\text{max}}[\boldsymbol{P}(t_{0})]\|\boldsymbol{\psi}(t_{0})\|_{2}^{2}.$ (10)
Thus, we have $\|\boldsymbol{\psi}(t)\|_{2}\leq\sqrt{\frac{\mu_{\text{max}}[\boldsymbol{P}(t_{0})]}{\mu_{\text{min}}[\boldsymbol{P}(t)]}}e^{\overline{\lambda}(t-t_{0})}\|\boldsymbol{\psi}(t_{0})\|_{2}\leq\sqrt{\frac{\mu_{\text{max}}[\boldsymbol{P}(t_{0})]}{\epsilon}}e^{\overline{\lambda}(t-t_{0})}\|\boldsymbol{\psi}(t_{0})\|_{2}$ using Eq. (7b). ∎
Theorem 1 will prove the stability if Eqs. (7b) and (7c) are feasible with $\overline{\lambda}=0$. It also allows us to analyze the growth rate for instabilities ($\overline{\lambda}>0$), which will be a main focus of this paper. We then discretize $\boldsymbol{P}(t)$ as $\boldsymbol{P}(t_{i})$ and discretize $\boldsymbol{A}(t)$ as $\boldsymbol{A}(t_{i})$, i=0,1,,ni=0,1,...,n with a fixed time step $\Delta t=t_{i+1}-t_{i}$. Because the time derivative cannot be directly implemented in our LMI solver, we use the forward Euler method to approximate it as $\dot{\boldsymbol{P}}(t_{i})=\frac{\boldsymbol{P}(t_{i+1})-\boldsymbol{P}(t_{i})}{\Delta t}$, leading to the discretized LMI as:
@@
$\dot{\boldsymbol{P}}(t_{i})+\boldsymbol{A}(t_{i})^{*}\boldsymbol{P}(t_{i})+\boldsymbol{P}(t_{i})\boldsymbol{A}(t_{i})-2\overline{\lambda}\boldsymbol{P}(t_{i})\preceq 0.$ (11c)
For a time-periodic system, we can enforce periodicity on $\boldsymbol{P}(t)$ such that $\boldsymbol{P}(t_{n})=\boldsymbol{P}(t_{0})$ [2, 22], and we only need to consider one period for the discretized LMI. Thus, we select $t_{0}=0$ and $t_{n}=T$, where $T=\frac{2\pi}{\omega}$ is the oscillation period. Although it is possible to first discretize the state-space model in Eq. (5) and then search for a Lyapunov function for the discrete-time system, here we formulate an LMI in Theorem 1 for a continuous-time system and then discretize the LMI, allowing us to define the upper bound of growth rate $\overline{\lambda}$ for a direct comparison with numerical simulations [17] and Floquet theory. Moreover, as this work considers a linear system in (5), the region of attraction will be the entire $\boldsymbol{\psi}\in\mathbb{C}^{5\times 1}$ for stable parameter regimes.
In Theorem 1, the $\boldsymbol{P}(t)\in\mathbb{C}^{5\times 5}$ matrix is complex because the state variables $\boldsymbol{\psi}\in\mathbb{C}^{5\times 1}$ are Fourier coefficients and in general complex. However, because $\boldsymbol{A}(t)\in\mathbb{R}^{5\times 5}$ here in (LABEL:eq:a_matrix) is a real matrix (which is not always the case for fluid systems, e.g., [21]), we can assume the weighting matrix $\boldsymbol{P}(t)\in\mathbb{R}^{5\times 5}$ and $\boldsymbol{P}(t_{i})\in\mathbb{R}^{5\times 5}$ as real symmetric matrices without loss of generality. The value of $\epsilon$ does not influence the upper bound of the growth rate $\overline{\lambda}$ due to the homogeneity of the constraints in (7c) and (11c), and thus, we set $\epsilon=1$ without loss of generality. We use YALMIP [14] to formulate our LMI in (11) and use MOSEK [15] as the semi-definite programming (SDP) solver. We first check the stability of the system by checking the feasibility of constraints in (11) with $\overline{\lambda}=0$, which is much quicker than calculating the value of $\overline{\lambda}$. If infeasible, we then solve (11) to minimize the upper bound of the growth rate $\overline{\lambda}>0$ through bisection search over $\overline{\lambda}$. All default settings are used for the YALMIP solver, where the bisection absolute gap tolerance of $10^{-5}$ is used.